a twin-engine airplane has to travel 896 miles from city a to city b. It takes 4 hours to get there with the wind. against the wind it takes 7 hours. what is the rate of the wind?
Answer by lwsshak3(11628)

a twin-engine airplane has to travel 896 miles from city a to city b. It takes 4 hours to get there with the wind. against the wind it takes 7 hours. what is the rate of the wind
**
let x=speed of airplane with no wind
let c=speed of wind
x+c=net speed of plane downwind
x-c=net speed of plane upwind
..
travel time =distance/speed
896/(x-c)=7
896/x+c)=4
..
7x-7c=896
4x+4c=896
..
28x-28c=3584
28x+28c=6272
subtract
56c=2688
c=48 mph